OpenStudy (anonymous):

A square is inscribed inside a circle. The square's vertices have coordinates R(0, 4), S(6, 2), T(4, -4), and U(-2, -2). Find the equation of the circle. (x - )^2 + y2 =

OpenStudy (anonymous):

Center of circle is center of square and radius is (side length)/sqrt(2)

OpenStudy (dumbcow):

 center = midpoint of RT diameter = length of RT center = (2,0) diameter = sqrt(4^2 + 8^2) = sqrt(80) = 4sqrt(5) --> radius = 2sqrt(5) circle: (x-2)^2 + y^2 = 20